Solve for x given: (5+i)x+(1-3i)=(2-3i)x+(2-i)
* ( i ) is understood to be the square root of -1

Answer by stanbon(48535)

(5+i)x+(1-3i)=(2-3i)x+(2-i)
Get the "x" terms together as follows:
(5+i)x-(2-3i)x=(2-i)-(1-3i)
[3+4i]x=(1+2i)
x=[(1+2i)(3-4i)]/[(3+4i)(3-4i)]
x=[3+8+6i-4i]/25
x=[11-2i]/25
Cheers,
Stan H.